# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v2

@@ 3.4. Quality Attributes and Tactics (RQ3)
Reliability: This QA is about the system’s correct operation and consistent performance without failure over a specified period, while ensuring the availability of the system. We identified XX studies that report on reliability, along with characteristics such as maturity, availability, and fault tolerance. We identified nine tactics overall (see Table 10 and the QA and Tactics Sheet in (Waseem et al., 2024)) to achieve reliability from the selected studies. These include deploying redundant engines for fault tolerance, distributing resources and replicating applications for improved response time, enforcing redundancy and distributed services for availability, deploying parallel search and multi-cloud distribution, and building fault-tolerant systems to enhance system reliability.
@@
Flexibility: This QA represents the system’s capacity to accommodate new features based on user requirements with minimal effort. We identified 10 studies (see Table 10 and the QA and Tactics Sheet in (Waseem et al., 2024)) that report this QA, along with five tactics that can help achieve it. For example, flexibility can be improved through tactics such as service-oriented architecture, on-demand dynamic allocation of resources across different cloud platforms, containers, container orchestration, model-driven development, risk analysis, and cross-cloud service orchestration. Using multi-cloud environment, for instance, offers the flexibility to choose services from different providers as per specific needs.
@@
[…]
Table 10. Quality attributes and tactics for containerized applications in multi-cloud environment
@@
Performance (Efficiency) Time behavior, Resource utilization, Capacity, Throughput, Response time (Almeida et al., 2014; Jha et al., 2019; Hua et al., 2020; Deng et al., 2022; Verdugo et al., 2017; Chen et al., 2022; Pham and Pham, 2015), (Ouyang et al., 2022; Ferry et al., 2014; Cui et al., 2016; Wang et al., 2020; Gao et al., 2017; Alonso et al., 2019; Gupta et al., 2021), (Rios et al., 2017; Carpio et al., 2020; Lorenzo et al., 2019; Dimitrijevic et al., 2019; Pandey et al., 2022; Altran et al., 2022; Shi et al., 2020), (Dreibholz et al., 2019; Wurster et al., 2018; Libardi et al., 2014; Harwalkar et al., 2019b; Casola et al., 2017; Bucur et al., 2018; Tricomi et al., 2017), (Sitaram et al., 2018; Rodigari et al., 2021; Janarthanan et al., 2018; Rahman and Lama, 2019; Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020) Efficient resource optimization and allocation
@@
Security Confidentiality, Authenticity, Access control, Authorization, Privacy, Trustworthiness, Integrity (Almeida et al., 2014; Gattobigio et al., 2022; Deng et al., 2022; Verdugo et al., 2017; Pham and Pham, 2015; Wang et al., 2020; Pahl, 2015), (Gao et al., 2017; Gupta et al., 2021; Rios et al., 2017; Dimitrijevic et al., 2019; Pandey et al., 2022; Bélair et al., 2019; Wurster et al., 2018), (Libardi et al., 2014; Casola et al., 2017; Bucur et al., 2018; Bolivar et al., 2018; Tricomi et al., 2017; Rodigari et al., 2021; Di Modica et al., 2019), (Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020; Kiss, 2018; Graupner et al., 2015a; Alaluna et al., 2017; Guechi and Maamri, [n. d.]; Ranjbar et al., 2017; Gundu et al., 2020) Encryption and strong authentication mechanisms
@@
Compatibility Interoperability, Co-existence (Pham and Pham, 2015; Ouyang et al., 2022; Cui et al., 2016; Pahl, 2015; Alonso et al., 2019; Baresi et al., 2016; Wurster et al., 2018), (Trihinas et al., 2015; Harwalkar et al., 2019b; Pellegrini et al., 2017; Harwalkar et al., 2019a; Graupner et al., 2015a; Ranjbar et al., 2017; Quint and Kratzke, 2018; Ferry et al., 2013) Implementing lightweight communication protocols and modes
@@
Scalability Capacity, Extensibility, Elasticity, Throughput, Responsiveness (Almeida et al., 2014; Pham et al., 2015; Verdugo et al., 2017; Chen et al., 2022; Pham and Pham, 2015; Ouyang et al., 2022; Slominski et al., 2015), (Pai and Kumar, 2021; Alhamazani et al., 2014; Cui et al., 2016; Wong et al., 2019; Pahl, 2015; Gao et al., 2017; Alonso et al., 2019), (Gupta et al., 2021; Mittal and Risco-Martín, 2017; Carpio et al., 2020; Shi et al., 2020; Baresi et al., 2016; Wurster et al., 2018; Trihinas et al., 2015), (Bucur et al., 2018; Tricomi et al., 2017; Rahman and Lama, 2019; Aral et al., 2020; Kiss, 2018; Alaluna et al., 2017; Mfula and Nurminen, 2018), (Pop et al., 2016; Ranjbar et al., 2017; Ferry et al., 2013; Kritikos et al., 2019)
@@
Reliability Maturity, Availability, Fault tolerance, Recoverability (Almeida et al., 2014; Jha et al., 2019; Ramos et al., 2021; Sabbioni et al., 2020; Naik, 2016a; Slominski et al., 2015; Alhamazani et al., 2014), (Ferry et al., 2014; Wong et al., 2019; Gao et al., 2017; Alonso et al., 2019; Al-shammari and Alwan, 2018; Sami and Mourad, 2020; Carpio et al., 2020), (Dimitrijevic et al., 2019; Wurster et al., 2018; Libardi et al., 2014; Buyya and Barreto, 2015; Acquaviva et al., 2017; Bucur et al., 2018; Bolivar et al., 2018), (del Castillo et al., 2013; Tricomi et al., 2017; Janarthanan et al., 2018; Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020; Tang, 2021), (Graupner et al., 2015a; Guechi and Maamri, [n. d.]; Mfula and Nurminen, 2018; Shen et al., 2017; Ferry et al., 2013) Deploying Redundant Engines for Fault-Tolerance
@@
Portability Installability, Replaceability, Adaptability (Pai and Kumar, 2021; Pahl, 2015; Gao et al., 2017; Mittal and Risco-Martín, 2017; Rahman and Lama, 2019; Ranjbar et al., 2017) Virtual Machine-Based Packaging
@@
Flexibility Functional completeness, Functional correctness, Functional appropriateness (Ouyang et al., 2022; Cui et al., 2016; Alonso et al., 2019; Mittal and Risco-Martín, 2017; Shi et al., 2020; Di Modica et al., 2019; Janarthanan et al., 2018), (Lazuka et al., 2022; Pop et al., 2016; Kritikos et al., 2019) Service-oriented Architecture
@@
Maintainability Modularity, Reusability, Analyzability, Modfiabality, Testability (Pham et al., 2015; Slominski et al., 2015; Gao et al., 2017; Alonso et al., 2019; Rahman and Lama, 2019; Mfula and Nurminen, 2018; Ferry et al., 2013) Microservices, easier maintenance
@@
Usability Appropriateness recognizability, Learnability, Operability , User interface aesthetics, Accessibility (Verdugo et al., 2017; Slominski et al., 2015; Alonso et al., 2019, 2019; Dimitrijevic et al., 2019; Mfula and Nurminen, 2018) Feedback-loop controller for multi-cloud Infrastructure
@@
3.5. Security Challenges and Solution Framework (RQ4 and RQ5)
@@
Data Security in Container-based Applications category covers multiple challenges related to data security in container-based applications, including data protection, database security, data compliance, and secure data transfer in a multi-cloud environment. Out of the 121 reviewed studies, 11 specifically highlighted these issues and provided corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The identified solutions focus on implementing robust encryption techniques, secure container orchestration, and compliance with regulatory standards such as GDPR and HIPAA. The solutions also include using Role-Based and Attribute-Based Access Controls (RBAC and ABAC) to manage secure access effectively. Standardization efforts like Secure CSP Interoperability and OpenStack API integration aim to mitigate risks of unauthorized access and vendor lock-in. Furthermore, comprehensive frameworks like DRA emphasize secure data management, ensuring privacy, fault tolerance, and secure data transfers across different cloud platforms.
Table 11. Identified security challenges and solutions for containerized applications in multi-cloud environments
@@
(Slominski et al., 2015) Data Protection Multi-Cloud Data Protection through Secure Containers
(Gupta et al., 2021) Database Security Enhanced multi-cloud Database Security via Containerization
(Dimitrijevic et al., 2019) Data Compliance Multi-cloud Legal Compliance and Data Protection
(Pandey et al., 2022) Data Security Secure CSP Interoperability and Selection in multi-cloud
(Acquaviva et al., 2017) Data Ownership and Privacy Multi-cloud Security via NoMISHAP Service Abstraction
(Sitaram et al., 2018) Data Transfer Security Secure multi-cloud Integration via OpenStack Standardization
(Bou Ghantous and Gill, 2021) Cloud Data Management Multi-cloud Data Security via DRA Framework
(Ahmad and Aujla, 2023) Data leaks and access risks Cryptography, smart contracts, encryption
(Sofia et al., 2024) Decentralized data security challenge Secure data handling with decentralized AI
(Neeraj et al., 2023) SLA data integrity challenge Blockchain for SLA security and integrity
(Santos et al., 2023) Multi-cloud data privacy issue Stringent security controls for data protection
@@
(Gattobigio et al., 2022) Connectivity and Access Control Multi-cloud Containerized Application Security
(Dreibholz et al., 2019) Secure Communication Middleware Secure Deployment for multi-cloud
(Buyya and Barreto, 2015) Provider-Agnostic Operations Aneka-Based Security and Performance for Containerized multi-cloud Applications
(Hossain et al., 2023) Microservices privacy and access control Encryption, authentication, service mesh for security
(Fugkeaw, 2023) Decentralized IAM security issue Blockchain for IAM security and data integrity
(Roy and Kalafatis, 2023) Secure communication across edge-cloud environments WireGuard for secure edge-cloud communication
(Benmerar et al., 2023) Multi-domain edge communication security Secure communication and AI anomaly detection
(Huang and Wang, 2023) Secure robot-cloud communication Secure communication and encryption for data integrity
@@
(Deng et al., 2022) Co-resident Security Secure Placement Strategy via Deep RL
(Gao et al., 2017) Container Isolation Power and Leakage Management for multi-cloud Security
(Bélair et al., 2019) Container and Host Security Host-Container Cooperation for multi-cloud Defense
(Wang et al., 2023) Container security and malware detection DockerWatch for malicious activity detection
(Tiwari and Sharma, 2023) Multi-tenant serverless security challenge Security best practices for serverless deployments
@@
(Ferry et al., 2014) Data Deployment Secure Private Data Deployment Strategy
(Wurster et al., 2018) Infrastructure Orchestration Standard Interface Implementation for Serverless Security
(Libardi et al., 2014) Storage Security Multi-cloud Storage Security with MSSF
(Quint and Kratzke, 2018) Migration Security Infrastructure-Aware and Agnostic Secure Migration
(Solayman and Qasha, 2023) Distributed infrastructure security challenge Orchestration security for IoT data protection
(Kataru et al., 2023) Securing AWS app deployments AWS best practices for CI/CD security
@@
(Sheridan et al., 2017) System Vulnerabilities Automated Security Evaluation and Bootstrapping
(Trihinas et al., 2015) Breach Detection and Prevention Metric Filtering and Fault Recovery for multi-cloud Security
(Lee and Nam, 2023) Network policy misconfiguration risks Automated network policy discovery and verification
(Nguyen and Chien, 2023) Stream processing security risks RBAM framework for secure data handling
@@
(Casola et al., 2017) Application Security and Compliance Risk Analysis and Security SLAs with MUSA Framework
(Harwalkar et al., 2019b) Data Security and Standardization OpenStack API-based multi-cloud Security and Interoperability
(Tricomi et al., 2017) Network Security Inter-cloud Communication Protection with SFC
(Wahab et al., 2016) Trust and Community Formation Trust-based Hedonic Coalitions for multi-cloud Security
(Kritikos et al., 2019) Orchestration and Adaptation Multi-cloud Service Orchestration and Cross-Level Adaptation
(Petrovski and Gusev, 2024) Multicloud compliance and security issue AWS security measures for multicloud environments
(Alonso et al., 2023a) Cross-cloud security and compliance Security frameworks for multi-cloud environment
@@
(Sami and Mourad, 2020) Placement Strategy Security Model for Volunteering Fog Services
(Pandi et al., 2023) Multi-cloud security concerns Security groups, VPCs, and infrastructure security
(Ramesh et al., 2023) Multicloud environment security issue Multicloud security for risk diversification
@@
(Sofia et al., 2023) Edge-cloud security concerns Enhanced security protocols for decentralized environments
@@
(Harzenetter et al., 2023) Cross-technology update security risks Integrated security in workflows for management
(Nagarajan et al., 2023) Resource allocation security risks Security measures in resource allocation framework
@@
Access and Communication Control category reports on multiple challenges related to maintaining secure multi-cloud connectivity and managing user access control. Out of the 121 reviewed studies, 9 specifically identified issues and proposed solutions addressing secure communication, provider-agnostic operations, and decentralized identity management (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). Advanced encryption and authentication mechanisms, e.g., TLS, JWT, and OAuth 2.0, were identified as solutions to secure communications across cloud and edge environments. Secure deployment of applications using middleware solutions was also discussed, which make use of SSL/TLS and access controls to enhance security attributes. Provider-agnostic middleware, e.g., the one based on Aneka, enables runtime provider selection and thus facilitates seamless operations across multi-cloud environment. Solutions for decentralized IAM security relate to blockchain utilization for improving data integrity and confidentiality. Secure communications in multi-domain edge environments are achieved through frameworks like WireGuard, backed by strong service mesh systems that maintain connectivity and protect against unauthorized access. Security monitoring for anomaly detection and policy enforcement was also noted as important for sustaining secure operations.
@@
Container Security category reports on challenges such as co-resident security, container isolation, host security, and multi-tenant serverless security in multi-cloud environment. Among the 121 reviewed studies, 5 identified these issues and proposed solutions to address them (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The solutions identified include the use of secure placement strategies via deep reinforcement learning to manage secure container placement in multi-cloud platforms. Security measures for container isolation include the technique of Power and Leakage Management, followed by Linux kernel-level isolation using SELinux. In order to enhance security on both container and host sides, solutions such as Host-Container Cooperation and Docker-based detection tools, like DockerWatch, enhance the defenses and malicious activity detection. Finally, with a view to further securing multi-tenant serverless environments, the implementation of best practices regarding security in the CLI tool used for managing serverless deployments was underlined.
@@
Infrastructure and Deployment Management Security category reports on challenges related to securing data deployment, infrastructure orchestration, storage security, and migration security in multi-cloud environment. Out of the reviewed 121 studies, 6 specifically addressed these types of challenges and proposed solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The identified solutions include the use of a Secure Private Data Deployment Strategy with encryption techniques like AES-256 and secure protocols such as SFTP and HTTPS to safeguard data. In the case of infrastructure orchestration, the focus is on the implementation of standard interfaces, such as OpenStack APIs, to ensure security and interoperability during deployments. For storage security, the solutions introduced the Multi-Cloud Storage Framework (MSSF), providing secure storage management through the implementation of RBAC and data encryption, along with periodic audits. The solutions also introduced measures for secure migration through both Infrastructure-Aware and Agnostic Secure Migration strategies, emphasizing image signing and data encryption under the least-privilege principle. Additionally, the reviewed studies proposed orchestration security for IoT distributed data and operations, as well as AWS best practices to help protect CI/CD pipelines running in multi-cloud environment.
@@
Security Monitoring and Breach Prevention category reports on challenges related to system vulnerabilities, breach detection and prevention, and network policy misconfiguration risks in multi-cloud environment. Out of the 121 reviewed studies, 4 specifically addressed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The identified solutions suggest implementing automated security evaluations and bootstrapping mechanisms for Virtual Machines with security configurations using tools like Terraform, Chef, Ansible, or Puppet. The studies also highlighted several other tools for vulnerability assessment, such as OpenVAS and Nessus, which can be used to address delays in patching and reduce system vulnerabilities. In terms of breach detection and prevention, the reviewed studies emphasized the need to enhance security through metric filtering, fault recovery, and the automated deployment of SIEM, IDS, and IPS tools. Log analysis solutions were considered necessary for proactive breach management. As a result, automated network policy discovery and verification mechanisms, such as KUNERVA, were proposed, incorporating rigorous verification to prevent most potential breaches in network policy management.
@@
Trust and Compliance Management category reports on solutions related to ensuring security and compliance in multi-cloud environment. Among the 121 reviewed studies, 7 specifically discussed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The identified solutions emphasize the need for security controls and standardized practices, such as risk analysis and secure coding using MUSA. The use of standardized API-based cloud interfaces, like the OpenStack API, was also emphasized to ensure both data security and interoperability between different cloud platforms. Establishing trust-based coalitions for community security was another important approach, supported by IAM systems and secure communication frameworks like SFC to enable secure interactions and collaborations within multi-cloud settings.

### v3

@@ 3.4 Quality Attributes and Tactics (RQ3)
Reliability: This QA is about the system’s correct operation and consistent performance without failure over a specified period, while ensuring the availability of the system. We identified 33 studies that report on reliability, along with characteristics such as maturity, availability, and fault tolerance. We identified nine tactics overall (see Table 10 and the QA and Tactics Sheet in [11]) to achieve reliability from the selected studies. These include deploying redundant engines for fault tolerance, distributing resources and replicating applications for improved response time, enforcing redundancy and distributed services for availability, deploying parallel search and multi-cloud distribution, and building fault-tolerant systems to enhance system reliability.
@@
Flexibility: This QA represents the system’s capacity to accommodate new features based on user requirements with minimal effort. We identified 10 studies (see Table 10 and the QA and Tactics Sheet in [11]) that report this QA, along with five tactics that can help achieve it. For example, flexibility can be improved through tactics such as service-oriented architecture, on-demand dynamic allocation of resources across different cloud platforms, containers, container orchestration, model-driven development, risk analysis, and cross-cloud service orchestration. Using multi-cloud environment, for instance, offers the flexibility to choose services from different providers as per specific needs.
@@
Takeaway 7: QAs and their associated tactics from the reviewed studies reveal that ensuring performance, security, reliability, and other QAs in containerized applications requires context-specific adaptations when operating in multi-cloud environments. Unlike single-cloud setups, these environments introduce heterogeneity in infrastructure, distributed control, and varying security models across providers. As a result, tactics such as SLA-aware resource optimization, federated access control, decentralized fault-tolerance mechanisms, and cross-platform provisioning are essential to achieving QA goals.
Table 10: Quality attributes and tactics for containerized applications in multi-cloud environment
@@
Performance (Efficiency) Time behavior, Resource utilization, Capacity, Throughput, Response time [95, 96, 93, 69, 120, 102, 80], [121, 89, 126, 66, 70, 105, 103], [56, 30, 113, 61, 67, 114, 97], [122, 140, 79, 78, 127, 133, 128], [81, 35, 123, 53, 108, 62, 48] Efficient resource optimization and allocation
@@
Security Confidentiality, Authenticity, Access control, Authorization, Privacy, Trustworthiness, Integrity [95, 34, 69, 120, 80, 66, 60], [70, 103, 56, 61, 67, 71, 140], [79, 127, 133, 40, 128, 35, 129], [108, 62, 48, 63, 57, 124, 58, 68, 64] Encryption and strong authentication mechanisms
@@
Compatibility Interoperability, Co-existence [80, 121, 126, 60, 105, 31, 140], [84, 78, 73, 108, 57, 68, 50, 51] Implementing lightweight communication protocols and modes
@@
Scalability Capacity, Extensibility, Elasticity, Throughput, Responsiveness [95, 86, 120, 102, 80, 121, 44], [59, 83, 126, 111, 60, 70, 105], [103, 106, 30, 97, 31, 140, 84], [133, 128, 53, 48, 63, 124, 42], [32, 68, 51, 76]
@@
Reliability Maturity, Availability, Fault tolerance, Recoverability [95, 96, 132, 65, 148, 44, 83], [89, 111, 70, 105, 112, 29, 30], [61, 140, 79, 98, 146, 133, 40], [107, 128, 123, 108, 62, 48, 41], [57, 58, 42, 109, 51] Deploying Redundant Engines for Fault-Tolerance
@@
Portability Installability, Replaceability, Adaptability [59, 60, 70, 106, 53, 68] Virtual Machine-Based Packaging
@@
Flexibility Functional completeness, Functional correctness, Functional appropriateness [121, 126, 105, 106, 97, 129, 123], [74, 32, 76] Service-oriented Architecture
@@
Maintainability Modularity, Reusability, Analyzability, Modfiabality, Testability [86, 44, 70, 105, 53, 42, 51] Microservices, easier maintenance
@@
Usability Appropriateness recognizability, Learnability, Operability , User interface aesthetics, Accessibility [120, 44, 105, 105, 61, 42] Feedback-loop controller for multi-cloud Infrastructure
@@
3.5 Security Challenges and Solution Framework (RQ4 and RQ5)
@@
Data Security in Container-based Applications category covers multiple challenges related to data security in container-based applications, including data protection, database security, data compliance, and secure data transfer in a multi-cloud environment. Out of the 121 reviewed studies, 11 specifically highlighted these issues and provided corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]). The identified solutions focus on implementing robust encryption techniques, secure container orchestration, and compliance with regulatory standards such as GDPR and HIPAA.
Table 11: Identified security challenges and solutions for containerized applications in multi-cloud environment
@@
[44] Data Protection Multi-Cloud Data Protection through Secure Containers
[103] Database Security Enhanced multi-cloud Database Security via Containerization
[61] Data Compliance Multi-cloud Legal Compliance and Data Protection
[67] Data Security Secure CSP Interoperability and Selection in multi-cloud
[146] Data Ownership and Privacy Multi-cloud Security via NoMISHAP Service Abstraction
[81] Data Transfer Security Secure multi-cloud Integration via OpenStack Standardization
[90] Cloud Data Management Multi-cloud Data Security via DRA Framework
[143] Data leaks and access risks Cryptography, smart contracts, encryption
[138] Decentralized data security challenge Secure data handling with decentralized AI
[147] SLA data integrity challenge Blockchain for SLA security and integrity
[125] Multi-cloud data privacy issue Stringent security controls for data protection
@@
[34] Connectivity and Access Control Multi-cloud Containerized Application Security
[122] Secure Communication Middleware Secure Deployment for multi-cloud
[98] Provider-Agnostic Operations Aneka-Based Security and Performance for Containerized multi-cloud Applications
[37] Microservices privacy and access control Encryption, authentication, service mesh for security
[144] Decentralized IAM security issue Blockchain for IAM security and data integrity
[46] Secure communication across edge-cloud environments WireGuard for secure edge-cloud communication
[104] Multi-domain edge communication security Secure communication and AI anomaly detection
[49] Secure robot-cloud communication Secure communication and encryption for data integrity
@@
[69] Co-resident Security Secure Placement Strategy via Deep RL
[70] Container Isolation Power and Leakage Management for multi-cloud Security
[71] Container and Host Security Host-Container Cooperation for multi-cloud Defense
[142] Container security and malware detection DockerWatch for malicious activity detection
[131] Multi-tenant serverless security challenge Security best practices for serverless deployments
@@
[89] Data Deployment Secure Private Data Deployment Strategy
[140] Infrastructure Orchestration Standard Interface Implementation for Serverless Security
[79] Storage Security Multi-cloud Storage Security with MSSF
[50] Migration Security Infrastructure-Aware and Agnostic Secure Migration
[38] Distributed infrastructure security challenge Orchestration security for IoT data protection
[134] Securing AWS app deployments AWS best practices for CI/CD security
@@
[55] System Vulnerabilities Automated Security Evaluation and Bootstrapping
[84] Breach Detection and Prevention Metric Filtering and Fault Recovery for multi-cloud Security
[145] Network policy misconfiguration risks Automated network policy discovery and verification
[119] Stream processing security risks RBAM framework for secure data handling
@@
[127] Application Security and Compliance Risk Analysis and Security SLAs with MUSA Framework
[78] Data Security and Standardization OpenStack API-based multi-cloud Security and Interoperability
[128] Network Security Inter-cloud Communication Protection with SFC
[72] Trust and Community Formation Trust-based Hedonic Coalitions for multi-cloud Security
[76] Orchestration and Adaptation Multi-cloud Service Orchestration and Cross-Level Adaptation
[100] Multicloud compliance and security issue AWS security measures for multicloud environments
[33] Cross-cloud security and compliance Security frameworks for multi-cloud environment
@@
[29] Placement Strategy Security Model for Volunteering Fog Services
[117] Multi-cloud security concerns Security groups, VPCs, and infrastructure security
[92] Multicloud environment security issue Multicloud security for risk diversification
@@
[118] Edge-cloud security concerns Enhanced security protocols for decentralized environments
@@
[141] Cross-technology update security risks Integrated security in workflows for management
[43] Resource allocation security risks Security measures in resource allocation framework
The solutions also include using Role-Based and Attribute-Based Access Controls (RBAC and ABAC) to manage secure access effectively. Standardization efforts like Secure CSP Interoperability and OpenStack API integration aim to mitigate risks of unauthorized access and vendor lock-in. Furthermore, comprehensive frameworks like DRA emphasize secure data management, ensuring privacy, fault tolerance, and secure data transfers across different cloud platforms.
@@
Access and Communication Control category reports on multiple challenges related to maintaining secure multi-cloud connectivity and managing user access control. Out of the 121 reviewed studies, 9 specifically identified issues and proposed solutions addressing secure communication, provider-agnostic operations, and decentralized identity management (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]). Advanced encryption and authentication mechanisms, e.g., TLS, JWT, and OAuth 2.0, were identified as solutions to secure communications across cloud and edge environments. Secure deployment of applications using middleware solutions was also discussed, which make use of SSL/TLS and access controls to enhance security attributes. Provider-agnostic middleware, e.g., the one based on Aneka, enables runtime provider selection and thus facilitates seamless operations across multi-cloud environment. Solutions for decentralized IAM security relate to blockchain utilization for improving data integrity and confidentiality. Secure communications in multi-domain edge environments are achieved through frameworks like WireGuard, backed by strong service mesh systems that maintain connectivity and protect against unauthorized access. Security monitoring for anomaly detection and policy enforcement was also noted as important for sustaining secure operations.
@@
Container Security category reports on challenges such as co-resident security, container isolation, host security, and multi-tenant serverless security in multi-cloud environment. Among the 121 reviewed studies, 5 identified these issues and proposed solutions to address them (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]). The solutions identified include the use of secure placement strategies via deep reinforcement learning to manage secure container placement in multi-cloud platforms. Security measures for container isolation include the technique of Power and Leakage Management, followed by Linux kernel-level isolation using SELinux. In order to enhance security on both container and host sides, solutions such as Host-Container Cooperation and Docker-based detection tools, like DockerWatch, enhance the defenses and malicious activity detection. Finally, with a view to further securing multi-tenant serverless environments, the implementation of best practices regarding security in the CLI tool used for managing serverless deployments was underlined.
@@
Infrastructure and Deployment Management Security category reports on challenges related to securing data deployment, infrastructure orchestration, storage security, and migration security in multi-cloud environment. Out of the reviewed 121 studies, 6 specifically addressed these types of challenges and proposed solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]). The identified solutions include the use of a Secure Private Data Deployment Strategy with encryption techniques like AES-256 and secure protocols such as SFTP and HTTPS to safeguard data. In the case of infrastructure orchestration, the focus is on the implementation of standard interfaces, such as OpenStack APIs, to ensure security and interoperability during deployments. For storage security, the solutions introduced the Multi-Cloud Storage Framework (MSSF), providing secure storage management through the implementation of RBAC and data encryption, along with periodic audits. The solutions also introduced measures for secure migration through both Infrastructure-Aware and Agnostic Secure Migration strategies, emphasizing image signing and data encryption under the least-privilege principle. Additionally, the reviewed studies proposed orchestration security for IoT distributed data and operations, as well as AWS best practices to help protect CI/CD pipelines running in multi-cloud environment.
@@
Security Monitoring and Breach Prevention category reports on challenges related to system vulnerabilities, breach detection and prevention, and network policy misconfiguration risks in multi-cloud environment. Out of the 121 reviewed studies, 4 specifically addressed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]). The identified solutions suggest implementing automated security evaluations and bootstrapping mechanisms for Virtual Machines with security configurations using tools like Terraform, Chef, Ansible, or Puppet. The studies also highlighted several other tools for vulnerability assessment, such as OpenVAS and Nessus, which can be used to address delays in patching and reduce system vulnerabilities. In terms of breach detection and prevention, the reviewed studies emphasized the need to enhance security through metric filtering, fault recovery, and the automated deployment of SIEM, IDS, and IPS tools. Log analysis solutions were considered necessary for proactive breach management. As a result, automated network policy discovery and verification mechanisms, such as KUNERVA, were proposed, incorporating rigorous verification to prevent most potential breaches in network policy management.
@@
Trust and Compliance Management category reports on solutions related to ensuring security and compliance in multi-cloud environment. Among the 121 reviewed studies, 7 specifically discussed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]). The identified solutions emphasize the need for security controls and standardized practices, such as risk analysis and secure coding using MUSA. The use of standardized API-based cloud interfaces, like the OpenStack API, was also emphasized to ensure both data security and interoperability between different cloud platforms. Establishing trust-based coalitions for community security was another important approach, supported by IAM systems and secure communication frameworks like SFC to enable secure interactions and collaborations within multi-cloud settings.